# Fairness for the People, by the People: Minority Collective Action

### v1

@@ 1 Introduction
As machine learning (ML) tools become increasingly accessible, more firms deploy them for decision-making. However, ML models often perpetuate societal biases present in their training data, leading to unfair outcomes across demographic groups [1]. Moreover, most fairness-preserving learning algorithms incur a non-negligible cost in accuracy or computational resources [2, 3, 4, 5], which can discourage practical adoption.
@@
Since firms control the training pipeline, end-users lack access to these algorithms and cannot directly enforce fair treatment. Yet, affected users routinely generate and share data — through clicks, ratings, or other contributions — that is used to train the firm’s models. Consequently, if underrepresented minority groups collaboratively alter the data they share, they might be able to steer the learned model towards fairer behavior, even without access to the firm’s training pipeline.
As an example, consider a multinational human resources company that makes a profit by filling vacancies and trains ML models on majority-derived hiring criteria. The company subcontracts recruiters from various communities including majority and minority (e.g. neighborhoods with a high concentration of an ethnic minority) to label the suitability of candidates for various job positions. A recruiter can choose to either use the firm’s template, such as a business-school degrees, or can substitute community-relevant signals such as informal apprenticeships. However, these community-relevant signals can sometimes contradict the company-suggested ones and moreover, unlike the company-suggested criteria, the local guidelines are not tested extensively and might not be as accurate.
@@
This idea is reminiscent of pre-processing fairness techniques [6, 7, 8, 9], which modify the data before model training. Unlike these prior approaches, which assume centralized control over the data, we consider the setting of algorithmic collective action [10, 11, 12, 13, 14], in which a small group of users strategically modifies their own data to influence the correlations learned by the model.
@@
We adapt the erasure strategy from Hardt et al. [10] to reduce predictive correlation between group membership and the target label by relabeling minority samples. The collective is restricted to members of the minority group since minority members are more motivated to join minority collective action [15, 16], and majority-group users may be less inclined to disrupt the status quo. We show that when a classifier is trained on data affected by this form of collective action, standard fairness metrics (e.g., demographic parity, equalized odds) improve substantially. This improvement is illustrated in Figure˜1, where a small collective of minority samples significantly reduces unfairness with minimal impact on prediction error.
@@
To summarize, our main contributions are: (1) We divert from the common firm-side fairness method and focus on user-side by introducing the setting of minority-only algorithmic collective action for fairness in ML (Section˜2), and design three algorithms that Pareto-dominate the random-choice baseline (Section˜3). (2) Through experiments on benchmark datasets, we demonstrate that these algorithms can significantly improve fairness metrics with only a slight accuracy penalty and few label flips, and minimal knowledge of majority-group data (Section˜4). (3) We investigate fundamental limitations of minority-only collectives and provide theoretical results showing that better representations and better counterfactual approximation methods can improve these algorithms (Section˜5).
@@ 2 Collective Action for Fairness
To establish the connection between fairness and collective action, Section˜2.1 first defines the problem setting and how unfairness can be measured. Then, Section˜2.2 describes the theoretical framework of collective action and how it can be utilized to mitigate bias. Finally, Section˜2.3 formally relates between collective action to group fairness metrics through counterfactual fairness.
2.1 Group fairness for classification.
@@
In the group-fairness paradigm, the sensitive attribute $a\in\{0,1\}$ partitions the data into subgroups, and fairness criteria seek to ensure similar outcomes across these groups. Common metrics include statistical parity (SP) [17, 18] and equalized odds (EqOd) [19]. In this work, we focus primarily on violations of EqOd, formally defined as
@@
which measures the differences between true positive and false positive rates. Section˜A.1 provides formal definitions and further discussion of these metrics.
ERM-trained models tend to achieve low predictive error, but this often comes at the cost of large fairness violations under SP and EqOd [2, 3, 20, 21]. Despite significant progress in fairness research, most solutions have traditionally focused on firm-side solutions: pre-processing the dataset, in-processing modifications to the training algorithm, or post-processing the classifier’s predictions. These approaches almost always incur increased error or additional pipeline complexity, discouraging firms to deploy them in practice.
While most prior work has focused on firm-side solutions, this work shifts the focus to user-side methods that do not require the firm’s participation. Since users generate the training data, they can collectively influence the learned model by strategically modifying their own behavior. For instance, consider a digital platform that recommends content to a user based on classifier predicting engagement labels $y_{i}\in\left\{\text{will engage},\text{ will not engage}\right\}$. The classifier, trained on historical user interactions, may unintentionally rely on group membership rather than individual preferences when making recommendations for minority members. In response, users can coordinate to alter their interaction patterns, such as clicking on or avoiding certain items. This collective action affects the dataset in a way that steers the learned classifier toward fairer outcomes, and is generally studied under the field of algorithmic collective action [10].
@@ 2.2 Algorithmic collective action
In social sciences, collective action refers to the coordinated efforts of individuals working together to pursue a shared goal [22, 23]. Hardt et al. [10] adapt this notion to machine learning, proposing that a group of users, termed a collective, can strategically modify their data to align the behavior of a trained classifier $h$ with the collective’s goals. In this formulation, the training distribution is a mixture distribution $\mathcal{D}\sim\mathbb{P}_{\alpha}=\alpha\mathbb{P}^{*}+(1-\alpha)\mathbb{P}_{0}$, where $\mathbb{P}^{*}$ and $\mathbb{P}_{0}$ are the collective and base distributions, and $\alpha\in\left[0,1\right]$ denotes the proportion of the population that belongs to the collective.
@@
When users have agency over the training data, one possible form of collective action for fairness is to modify their features to increase correlation with the label $y=1$. An analogous firm-side approach is fair representation learning (FRL), which learns a transformation from the input space to a representation space such that ERM leads to a classifier that is both accurate and fair [8, 24]. However, a hindrance of FRL in the context of collective action is that the transformation must be applied consistently at inference time, requiring active cooperation from each minority member to transform their features. In contrast, our setting assumes users have control only over the labels and cannot intervene in other parts of the machine learning pipeline.
@@
To achieve signal erasure, Hardt et al. [10] propose the collective relabels itself with the most likely label under the transformation $g$. Formally, the strategy is defined as
$\left(x,y\right)\rightarrow\left(x,\mathop{\mathrm{argmax}}_{y^{\prime}\in\left\{0,1\right\}}\mathbb{P}_{0}\left(y^{\prime}|g\left(x\right)\right)\right).$ (4)
Since this strategy leaves the features unchanged, it is well-suited for settings where the minority is limited to modify only their labels, such as ours. For $\epsilon{}$-optimal Bayes classifiers (Definition˜2 in Section˜A.2), Hardt et al. [10] prove the following lower bound for its success
@@
where $\tau=\operatorname*{\mathbb{E}}_{x\sim\mathbb{P}_{0}}\left[\max_{y^{\prime}\in\left\{0,1\right\}}\left|\mathbb{P}_{0}\left(y^{\prime}|x\right)-\mathbb{P}_{0}\left(y^{\prime}|g\left(x\right)\right)\right|\right]$ captures the sensitivity of the true label distribution to the transformation $g$.
@@ 2.3 Counterfactual fairness
The concept of counterfactual fairness (CF) [25, 26, 27] bridges between signal erasure success to group fairness. To introduce this idea, assume that a sample $x$ is generated by a causal model, in which the group membership $A$ is a causal parent. Then a classifier $h$ is counterfactually fair if its predictions are invariant to interventions on the group membership, i.e., $h\left(x\right)=h\left(x_{A\leftarrow a^{\prime}}\right)$ for any $a^{\prime}$, where $x_{U\leftarrow u}$ denotes an intervention on a causal parent $U$ of a sample $x$. In certain causal contexts, CF implies or aligns with group fairness criteria such as SP or EqOd [28]. Therefore, if collective action induces a counterfactually fair classifier, it may also induce a fair classifier under SP or EqOd.
Since our focus is on fairness for the minority group, we relax the original definition of CF [25].
@@
By this definitions, changing the group membership of a minority individual, in a counterfactual sense, has no effect on the classifier’s prediction. Collective action can theoretically enforce such fairness by applying the erasure strategy from Equation˜4 with the counterfactual signal $g\left(x\right)=x_{A\leftarrow 0}$, which replaces a minority individual with its majority-group counterfactual. This collective action aligns the signal erasure success from Equation˜3 with minority-focused counterfactual fairness from Definition˜1. The following proposition, proved in Section˜B.1, formalizes this alignment.
@@
This result directly connects between collective action theory to fairness. Thus, perfect success of the collective is equivalent to achieving minority-focused counterfactual fairness.
(a) RB-label
(b) RB-dist
@@ 3 Approximating the Counterfactual Label
This section describes how a minority collective can approximate a signal-erasure strategy to promote fairness in practice. While the theory of signal erasure has been studied before [10, 14], prior work lacks empirical evaluation. In this paper, we present the first practical algorithm for signal erasure and provide experimental results in Section˜4. As discussed in Section˜2.3, a suitable signal to erase is $g\left(x\right)=x_{A\leftarrow 0}$, where each collective member relabels themselves according to Equation˜4.
However, end-users lack access to the true causal model and cannot compute the counterfactual labels directly. To address this limitation, we propose to assign each collective member $i$ a score $s_{i}$, which serves as a proxy for the likelihood that they would receive the label $y=1$ if they belonged to the majority. Given a budget of $M$ label flips, the collective selects the $M$ members with the highest scores; these individuals flip their labels from $y=0$ to $y=1$. The budget $M$ controls the accuracy–fairness tradeoff, where a higher budget typically leads to better fairness, but higher error (Figures˜5, 7 and 8).
We introduce three model-agnostic scoring functions, each capturing a different notion of similarity to majority users:
@@ 4 Experimental Results
This section evaluates the performance of our proposed methods under varying settings. We compare the three methods, RB-label, RB-dist, RB-prob, against a random baseline that flips $y=0$ labels to $y=1$ for $M$ randomly selected collective members. We conducted experiments on the tabular datasets COMPAS [29], Adult [30], HSLS [31], ACS-Income [32], the image dataset Waterbirds [33] and the text dataset CivilComments [34]. For Waterbirds, we use features extracted from a pre-trained ResNet-18 (denoted Waterbirds-Full) and for CivilComments, we used the extracted features from Hugging Face’s pre-trained bert-base-uncased model (denoted CivilComments-Full). In addition to the complete features of Waterbirds and CivilComments, we also include experiments on the PCA features, with 85 components for Waterbirds (denoted Waterbirds-PCA) and 100 components for CivilComments (denoted CivilComments-PCA). Details on the datasets and the pre-processing are provided in Section˜C.1.
All reported metrics are computed on a fixed test set, without any collective action, and averaged over 10 independent runs. In each run, we randomly selected a minority collective, which then applies one of the methods described in Section˜3. For the KNN-based methods, we tuned the neighborhood size $k$ using a $15\%$ validation split from the train set, optimizing for EqOd and SP. Finally, we trained a gradient-boosted decision tree on each modified train set. More technical details are given at Section˜C.2 and the complete set of results can be found in Section˜D.2.
@@
Flipping cost
Since each method scores candidates differently, they may also vary in efficiency, that is, the number of label flips required to achieve a given level of fairness. To evaluate efficiency, Figure˜4 plots EqOd as a function of number of label flips $M$, where lower curves indicate more efficient methods. The random baseline consistently yields the worst EqOd across all values of $M$, highlighting the value of informed relabeling algorithms. However, no single method dominates the others in all settings: While RB-prob and RB-label often outperform the other methods, RB-dist can surpass them in specific cases (e.g., Figure˜4(a)), or perform comparably to the random baseline in others (Figure˜4(c)).
These results suggest that a well-chosen scoring function enables the collective to achieve a desired level of fairness with fewer label flips, reducing the “cost” of collective action and mitigating the accuracy loss from excessive relabeling.
Interestingly, beyond a certain number of flips, EqOd begins to increase, indicating that excessive flipping can shift unfairness from the minority to the majority. This upturn reflects the fundamental limits of minority collective action for fairness, a point we elaborate on in Section˜5.

### v2

@@ 1 Introduction
As machine learning (ML) tools become increasingly accessible, more firms deploy them for decision-making. However, ML models often perpetuate societal biases present in their training data, leading to unfair outcomes across demographic groups [barocasBigDatasDisparate2016]. Moreover, most fairness-preserving learning algorithms incur a non-negligible cost in accuracy or computational resources [menonCostFairnessBinary2018, zhaoInherentTradeoffsLearning2019, dehdashtianUtilityfairnessTradeoffsHow2024, sadeghiCharacterizingTradeoffInvariant2022], which can discourage practical adoption.
@@
Since firms control the training pipeline, end-users lack access to these algorithms and cannot directly enforce fair treatment. Yet, affected users routinely generate and share data, through clicks, ratings, or other contributions, that is used to train the firm’s models. Consequently, if underrepresented minority groups collaboratively alter the data they share, they might be able to steer the learned model towards fairer behavior, even without access to the firm’s training pipeline.
For example, consider a human resources company that makes a profit by filling vacancies and trains ML models on resumes to predict the skills of the candidates. While the majority may have more formal education and college degrees, disadvantaged groups may have informal training or internships. As a result, the ML model does not assign the correct skills to the minority due to their lack of formal education, despite their practical experience. The minority members can react to this injustice by collectively submitting their resumes, but reframing their reported skills, such as sales or management. In Appendix˜A we describe the settings where such collective action is applicable and list other real-world examples.
@@
This idea is reminiscent of pre-processing fairness techniques [kamiranClassifyingDiscriminating2009, luongKNNImplementationSituation2011, zemelLearningFairRepresentations2013, madrasLearningAdversariallyFair2018], which modify the data before model training. Unlike these prior approaches, which assume centralized control over the data, we consider the setting of algorithmic collective action (ACA) [hardtAlgorithmicCollectiveAction2023, ben-dovRoleLearningAlgorithms2024, baumannAlgorithmicCollectiveAction2024, siggDeclineNowCombinatorial2025, gauthierStatisticalCollusionCollectives2025], in which a small group of users strategically modifies their own data to influence the correlations learned by the model.
@@
We adapt the erasure strategy from hardtAlgorithmicCollectiveAction2023 to reduce predictive correlation between group membership and the target label by relabeling minority samples. The collective is restricted to members of the minority group since minority members are more motivated to join minority collective action [saleemWhenHowNegative2021, begenyPowerIngroupPromoting2022], and majority-group users may be less inclined to disrupt the status quo. We show that when a classifier is trained on data affected by this form of ACA, standard fairness metrics (e.g., demographic parity, equalized odds) improve substantially. This improvement is illustrated in Figure˜1, where a small collective of minority samples significantly reduces unfairness with minimal impact on prediction error.
@@
To summarize, our main contributions are: (1) We divert from the common firm-side fairness methods and focus on user-side by introducing the setting of minority-only algorithmic collective action for fairness in ML (Section˜2), and design three algorithms that Pareto-dominate the random-choice baseline. (2) Through experiments on benchmark datasets, we demonstrate that these algorithms can significantly improve fairness metrics with only a slight accuracy cost and few label flips, and minimal knowledge of majority-group data. (3) We investigate fundamental limitations of minority-only collectives and provide theoretical results showing that better representations and better counterfactual approximation methods can improve these algorithms.
@@ 2 Collective Action for Fairness
To establish the connection between fairness and ACA, Section˜2.1 first defines the problem setting and how unfairness can be measured. Then, Section˜2.2 describes the theoretical framework of ACA and how it can be utilized to mitigate bias. Finally, Section˜2.3 formally relates between ACA to group fairness metrics through counterfactual fairness.
2.1 Group fairness for classification
@@
In the group-fairness paradigm, the sensitive attribute $a\in\{0,1\}$ partitions the data into subgroups, and fairness criteria seek to ensure similar outcomes across these groups. Common metrics include statistical parity (SP) [caldersBuildingClassifiersIndependency2009, dworkFairnessAwareness2012] and equalized odds (EqOd) [hardtEqualityOpportunitySupervised2016]. In this work, we focus primarily on violations of EqOd, formally defined as
@@
which measures the differences between true positive and false positive rates. Section˜B.1 provides formal definitions and further discussion of these metrics.
ERM-trained models tend to achieve low predictive error, but this often comes at the cost of fairness violations under SP and EqOd [menonCostFairnessBinary2018, zhaoInherentTradeoffsLearning2019, bardenhagenBoostingWorstgroupAccuracy2021, sanyalHowUnfairPrivate2022]. Despite significant progress in fairness research, most solutions have traditionally focused on firm-side solutions: pre-processing the dataset, in-processing modifications to the training algorithm, or post-processing the classifier’s predictions. These approaches almost always incur errors or additional pipeline complexity, discouraging firms to deploy them in practice.
While most prior work has focused on firm-side solutions, this work shifts the focus to user-side methods that do not require the firm’s participation. Since users generate the training data, they can collectively influence the learned model by strategically modifying their own behavior. For instance, consider a digital platform that recommends content to a user based on classifier predicting engagement labels $y_{i}\in\left\{\text{will engage},\text{ will not engage}\right\}$. The classifier, trained on historical user interactions, may unintentionally rely on group membership rather than individual preferences when making recommendations for minority members. In response, users can coordinate to alter their interaction patterns, such as clicking on or avoiding certain items. This ACA affects the dataset in a way that steers the learned classifier toward fairer outcomes, and is generally studied under the field of algorithmic collective action [hardtAlgorithmicCollectiveAction2023].
@@ 2.2 Algorithmic collective action
In social sciences, collective action refers to the coordinated efforts of individuals working together to pursue a shared goal [olsonCollectiveAction1989, marwellCriticalMassCollective1993]. hardtAlgorithmicCollectiveAction2023 adapt this notion to ML, proposing that a group of users, termed a collective, can strategically modify their data to align the behavior of a trained classifier $h$ with the collective’s goals. In this formulation, the training distribution is a mixture distribution $\mathcal{D}\sim\mathbb{P}_{\alpha}=\alpha\mathbb{P}^{*}+(1-\alpha)\mathbb{P}_{0}$, where $\mathbb{P}^{*}$ and $\mathbb{P}_{0}$ are the collective and base distributions, and $\alpha\in\left[0,1\right]$ denotes the proportion of the collective.
@@
With user agency over the data, one possible form of ACA for fairness is to modify their features to increase correlation with the label $y=1$. An analogous firm-side approach is fair representation learning (FRL), which learns a transformation from the input space to a representation space such that ERM leads to a classifier that is both accurate and fair [zemelLearningFairRepresentations2013, jovanovicFAREProvablyFair2023]. However, a hindrance of FRL in the context of ACA is that the transformation must be applied consistently at inference time, requiring active cooperation from each minority member to transform their features. In contrast, our setting assumes users have control only over the labels and cannot intervene in other parts of the machine learning pipeline.
@@
To achieve signal erasure, hardtAlgorithmicCollectiveAction2023 propose the collective relabels itself with the most likely label under the transformation $g$. Formally, the strategy is defined as
$x,y\rightarrow x,\mathop{\mathrm{argmax}}_{y^{\prime}\in\left\{0,1\right\}}\mathbb{P}_{0}\left(y^{\prime}|g\left(x\right)\right).$ (4)
Since this strategy leaves the features unchanged, it is well-suited for settings where the minority is limited to modify only their labels, such as ours. For $\epsilon{}$-optimal Bayes classifiers (Definition˜2 in Section˜B.2), hardtAlgorithmicCollectiveAction2023 prove the following lower bound for its success
@@
where $\tau=\operatorname*{\mathbb{E}}_{x\sim\mathbb{P}_{0}}\left[\max_{y^{\prime}\in\left\{0,1\right\}}\left|\mathbb{P}_{0}\left(y^{\prime}|x\right)-\mathbb{P}_{0}\left(y^{\prime}|g\left(x\right)\right)\right|\right]$ captures the sensitivity of $y$ under $g$.
@@ 2.3 Counterfactual fairness
The concept of counterfactual fairness (CF) [kusnerCounterfactualFairness2017, gargCounterfactualFairnessText2019, wuCounterfactualFairnessUnidentification2019] bridges between signal erasure success to group fairness. To introduce this idea, assume that a sample $x$ is generated by a causal model, in which the group membership $A$ is a causal parent. Then a classifier $h$ is counterfactually fair if its predictions are invariant to interventions on the group membership, i.e., $h\left(x\right)=h\left(x_{A\leftarrow a^{\prime}}\right)$ for any $a^{\prime}$, where $x_{U\leftarrow u}$ denotes an intervention on a causal parent $U$ of a sample $x$. In certain causal contexts, CF implies or aligns with group fairness criteria such as SP or EqOd [anthisCausalContextConnects2023]. Therefore, if ACA induces a counterfactually fair classifier, it may also induce a fair classifier under SP or EqOd.
As focus is on fairness for the minority, we relax the original definition of CF [kusnerCounterfactualFairness2017].
@@
By this definitions, changing the group membership of a minority individual, in a counterfactual sense, has no effect on the classifier’s prediction. ACA can theoretically enforce such fairness by applying the erasure strategy from Equation˜4 with the counterfactual signal $g\left(x\right)=x_{A\leftarrow 0}$, which replaces a minority individual with its majority-group counterfactual. This ACA aligns the signal erasure success from Equation˜3 with minority-focused counterfactual fairness from Definition˜1. The following proposition, proved in Section˜C.1, formalizes this alignment.
@@
This result directly connects between ACA theory to fairness. Thus, perfect success of the collective is equivalent to achieving minority-focused counterfactual fairness.
@@ 3 Approximating the Counterfactual Label
This section describes how a minority collective can approximate a signal-erasure strategy to promote fairness in practice. While the theory of signal erasure has been studied before [hardtAlgorithmicCollectiveAction2023, gauthierStatisticalCollusionCollectives2025], prior work lacks empirical evaluation. In this paper, we present the first practical algorithm for signal erasure and provide experimental results in Section˜4. As discussed in Section˜2.3, a suitable signal to erase is $g\left(x\right)=x_{A\leftarrow 0}$, where each collective member relabels themselves according to Equation˜4.
However, end-users lack access to the true causal model and cannot compute the counterfactual labels directly. To address this limitation, we propose to assign each collective member $i$ a score $s_{i}$, which serves as a proxy for the likelihood that they would receive the label $y=1$ if they belonged to the majority. Given a budget of $M$ label flips, the collective selects the $M$ members with the highest scores; these individuals flip their labels from $y=0$ to $y=1$. The budget $M$ controls the accuracy–fairness tradeoff, where a higher budget typically leads to better fairness, but higher error.
We introduce three scoring functions, each capturing a different notion of similarity to majority users:
@@ 4 Experimental Results
This section evaluates the performance of our methods. We compare the three methods, RB-label, RB-dist, RB-prob, against a random baseline that flips $y=0$ labels to $y=1$ for $M$ randomly selected collective members. We conducted experiments on the tabular datasets COMPAS [mattuHowWeAnalyzed2016], Adult [beckerAdult1996], HSLS [jeongFairnessImputationDecision2022], ACS-Income [dingRetiringAdultNew2021], the image dataset Waterbirds [sagawaDistributionallyRobustNeural2020] and the text dataset CivilComments [borkanNuancedMetricsMeasuring2019]. For Waterbirds, we use features extracted from a pre-trained ResNet-18 (denoted Waterbirds-Full) and for CivilComments, we used the extracted features from Hugging Face’s pre-trained bert-base-uncased model (denoted CivilComments-Full). In addition to the complete features of Waterbirds and CivilComments, we also include experiments on the PCA features, with 85 components for Waterbirds (denoted Waterbirds-PCA) and 100 components for CivilComments (denoted CivilComments-PCA). Details on the datasets and the pre-processing are provided in Section˜D.1.
All reported metrics are computed on a fixed test set, without any ACA, and averaged over 10 independent runs for each method described in Section˜3. In each run, we randomly selected a minority collective to apply the method. For the KNN-based methods, we tuned the neighborhood size $k$ using a $15\%$ validation split from the train set, optimizing for EqOd and SP. Finally, we trained a gradient-boosted decision tree on each modified train set. More technical details are given at Section˜D.2 and the complete set of results can be found in Section˜E.2.
@@
Flipping cost
Since each method scores candidates differently, they may also vary in efficiency, that is, the number of label flips required to achieve a given level of fairness. To evaluate efficiency, Figure˜4 plots EqOd as a function of number of label flips $M$, where lower curves indicate more efficient methods. The random baseline consistently yields the worst EqOd across all values of $M$, highlighting the value of informed relabeling algorithms. However, no single method dominates the others in all settings: While RB-prob and RB-label often outperform the other methods, RB-dist can surpass them in specific cases (e.g., Figure˜4(a)), or perform comparably to the random baseline in others (Figure˜4(c)). These results suggest that a well-chosen scoring function enables the collective to achieve a desired level of fairness with fewer label flips, reducing the “cost” of ACA and mitigating the accuracy loss from excessive relabeling.
Interestingly, beyond a certain number of flips, EqOd begins to increase, indicating that excessive flipping can shift unfairness from the minority to the majority. This upturn reflects the fundamental limits of minority ACA for fairness, a point we elaborate on in Section˜5.